### v3

# A Distributed Privacy Preserving Model for the Detection of Alzheimer’s Disease

Abstract
BACKGROUND: Segmentation of medical data, concerns about personal health information (PHI) breaches, and the direct and indirect costs of consolidating and managing such segmented date should motivate diagnostic machine learning (DML) researchers to identify privacy-preserving machine learning algorithms that can train on distributed or decentralized datasets of different modalities. Federated learning models provide such a decentralized machine learning framework in which multiple investigators in possession of disparate datasets and working on different devices or servers can train collaboratively a global machine learning models without ever having to exchange local data and thus can meet statutory PHI protections. To this end, a vertical federate learning model is devised and tested for efficacy in the detection of Alzheimer’s Disease (AD).
METHODS: The second version of Open Access Series of Imaging Studies– with its panoply of demographic, imaging, and clinical assessment datasets– was used to test a multimodal vertical federated learning (VFL) model for AD detection.
RESULTS: By training and validating this VFL model on the demographic, clinical, and MRI data in OASIS-2, an 82.9% accuracy rate is achieved, consistent with previously reported results.
CONCLUSIONS: The VFL architecture proposed herein offers a novel distributed architecture, enabling collaborative learning across diverse sources of medical data while respecting statutory privacy constraints. By leveraging multiple modalities of data, the robustness and accuracy of AD detection can be enhanced. This model not only contributes to the advancement of federated learning techniques but also holds promise for overcoming the hurdles posed by data segmentation in medical research.
1 Introduction
The increasing volume of medical data should be a priceless asset for advancing diagnostic machine learning (DML) models in healthcare 1; 2; 3. Commonly, these large datasets are consolidated onto a centralized server for collaborators to investigate. Unfortunately, consolidating medical data in such a fashion in order to train a centralized diagnostic machine learning model can be burdensome, costly, and impractical. For example, the Health Insurance Portability and Accountability Act (HIPAA) 4; 5 and as updated through the Health Information Technology for Economic and Clinical Health Act (HITECH)6 mandate safeguards to protect any personal health information (PHI) from any breaches. Additionally, consolidating medical data in order to train a centralized model often can be an impractical endeavor. The segmentation inherent in medical data across diverse sources imposes significant limitations on both the depth and breadth of available information for robust DML training 7. Not surprisingly, consolidating and managing this data is fraught with administrative burdens, technical challenges, and considerable expenses. Specifically, regional health information organizations (RHIOs) require US \$12 million each for development and incur \$2-3 million in operating costs 8; 9. Just as a physician must grasp different types of clinical data (such as a history, physical examination, laboratory results, imaging studies, and questionnaires) to develop a unifying diagnosis, each data source used to train a DML model provides only a partial perspective. DML demands novel methodologies capable of training from different modes of data while optimizing model development in any distributed environment. Thus, the utility of multimodal models becomes clearly apparent.
Arguably, as explained above, prevailing approaches to multimodal model training practice data consolidation 10 –fundamentally misaligned with the objectives of (1) preserving data segmentation, (2) protecting PHI, and (3) still optimizing model performance. On the other hand, the integration of diverse data modalities holds promise for a more comprehensive understanding, particularly in addressing complex medical conditions. Neuropsychiatric disorders, such as Alzheimer’s Disease (AD), require interpreting a host of different objective and subjective data to establish a definitive diagnosis. Because DML research on AD using multimodal data already has been reported, I chose to train a novel machine learning model– specifically, a vertical federated learning model– to detect AD. The decision to address AD not only stems from a technical standpoint but also aligns with a broader, urgent societal need. Specifically, in 2024, an estimated 6.9 million or 1 in 9 (11%) of all Americans aged 65 years and older are living with AD11. The Centers for Medicare and Medicaid Services shall pay out US\$ 206 billion for direct AD care12. Highlighting the overall the overall economic and associated burdens of AD to patients, their families, and society is the fact that there is an additionally estimated 18 billion unpaid hours for dementia-related care in 2022, valued at \$339.5 billion 13. The number of Americans aged 65 and older not only will increase exponentially in the United States but also elsewhere– as will the cost and other burdens of caring for AD. Early detection and intervention of AD should improve the care and cut the cost of care for AD. Therefore, there is a critical need for technical innovations facilitating early intervention and personalized care.
In this paper, a novel multimodal vertical federated learning (VFL) framework that can be used for AD diagnosis is proposed and implemented. First, an overview review of both horizontal and vertical federated learning initiatives is provided, along with previous research in DML in AD. Next, the multimodal dataset, which consists of both demographic data and MRI images, is introduced. I then detail the experimental setup and model architectures, which can be readily reproduced by other investigators. After discussing the results of this VFL for AD detection, I suggest how this framework can be improved and expanded so that it readily can be used in the clinical setting.
2 Background
Figure 1: A Horizontal Federated Learning architecture for three radiology departments in different institutions to collaboratively train a single model using all of their data without medical record exchange.
2.1 What Is Federated Learning?
Federated learning is a decentralized machine learning framework which allows multiple disparate parties– often on separate devices or servers– to collaboratively train a global machine learning model without ever having to exchange their local data 14. In scenarios wherever data security, patient privacy, and maintenance of data locality are necessary, federated learning can be a powerful tool. Rather than consolidating data in a central repository, federated learning leverages a federated architecture, where each data source (a ”client” or ”federate”) maintains its own dataset and participates in model training through local computations. The global model then is updated iteratively, aggregating the model updates generated by each client while preserving the privacy and security of individual data sources. Based on the framework architecture, federated learning models can be described as horizontal or vertical.
2.2 Horizontal Federated Learning
Horizontal Federated Learning (HFL) is defined as a scenario in which different parties share the same feature space but contain different data samples 15. After initialization of a global model, each client, possessing the same type of data, trains the model locally. Clients collaborate to improve the global model by sharing insights from their locally collected data with the central server. The central server then aggregates the models received from the clients, typically by averaging the model parameters. This averaging process ensures that the global model benefits from the diverse knowledge learned on different clients while preserving privacy. One common method used in this process is Federated Averaging (FedAvg) 14, where the central server orchestrates the federated learning process by specifically averaging the model updates and then redistributing the improved model to all clients. Therefore, HFL is effective in scenarios where multiple federates are collecting similar data samples from multiple federates (or entities) without having to consolidate this data onto a centralized server. In a clinical setting, one could imagine a scenario where there are three hospitals. These hospitals all contain the same fields of data (such as magnetic resonance imaging or MRI). Radiology Department A at one hospital may have MRIs for Patients 1-3 who were being evaluated for AD, Radiology Department B may MRIs of Patients 4-6, and Radiology Department C for Patients 7-9. In this case, an HFL model could be used for DML training since one mode of data (MRIs) from different patients are obtained from multiple federates in order to train a single neural network. Local proprietary interests of the data and PHI protections thereby would be preserved since the actual images and patient identification is neither stored centrally nor shared among the collaborators . The architecture of HFL is illustrated in figure 1.
2.3 Clinical Applications for Horizontal Federated Learning
There are many survey papers that both propose and implement HFL in the healthcare research 16; 17. In terms of primary research of HFL models for AD detection training, Khalil and colleagues 18 introduce a hardware accelerated horizontal federated multi-layer perceptron on field programmable gate arrays with 89% accuracy on blood plasma data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) 19. They also used K-nearest neighbors (KNN), support vector machines, decision trees, and logistic regression analyses.
Similarly, MRIs from the ADNI and Australian Imaging, Biomarkers and Lifestyle (AIBL) 20 datasets were utilized to implement a federated domain adaptation framework via Transformer (FedDAvT)21. This approach performed Unsupervised Deep Domain Adaptation and managed to achieve the following accuracy rates of: (1) 88.75% for the AD vs. Normal Controls (NC), (2) 69.51% for Mild Cognitive Impairment (MCI) vs. NC, and (3) 69.88% for AD vs. MCI binary classification tasks.
A process called federated conditional mutual learning (FedCM)–which considers the similarity between clients– improved these DML models that trained on MRI imaging 22. Using MRIs from the first version of Open Access Series of Imaging Studies (OASIS 1) 23, ADNI, and AIBL datasets and then by extending upon FedAvg through FedCM, accuracy rates of 75.5% accuracy rate on OASIS-1 and of 76.0% on AIBL and ADNI were achieved.
In each of the above studies of HFL models for AD detection, only one modality of data was used– whether blood plasma protein levels or MRIs. Nevertheless, some investigators have attempted to train HFLs on multimodal data. For example, Gao and colleagues performed a manifold projection of each electroencephalogram (EEG) signal, first by calculating a domain loss on each disparate signal and then by applying HFL to classify each signal24. Although only one type of testing was analyzed (EEGs), manifold projection is necessary since each EEG signal comes from a different area of the brain and thus can be considered multimodal data. Moreover, this paper confirmed that all user data privacy was preserved. Additionally, Peng and colleagues used federated graph convolution so that both imaging data and demographic data are converted into uniform high level features. A special graph generative adversarial network is then used to fill out missing nodes in the graph. Therefore, the multimodal data is projected onto a uniform feature space, where FedAVG is applied to the locally trained graph for detecting both Alzheimer’s Disease (AD) and Autism Spectrum Disorder (ASD) 25. Both of the technologies described above succeeded in their objectives, training either a single federated neural network using an HFL process which would not perform well without the manifold projection 24 or on multimodal data by using the same manifold projection25. Unfortunately, projecting features onto a different latent space inevitably results in loss of information. Additionally, by losing the structure of the data, it is not possible to apply algorithms that best suit the data (e.g. Convolutional Neural Networks on images). Finally, the manifold projection of certain datasets is not possible because it is extremely difficult to project images down to the same latent space as demographic features, for example.
2.4 Vertical Federated Learning
Vertical Federated Learning (VFL)– in contrast to HFL– addresses instances in which multiple federates share the feature space and each training example is spread across different entities 26. Clients possess different features of a common entity, and the goal is to train a model that leverages all features for improved prediction and analysis. This approach enables predictive modeling and analytics across heterogeneous, multimodal data sources while preserving data privacy and security 27.
In the clinical setting, for example, if a primary care physician (PCP) is concerned that a patient has early signs of AD, the PCP might order a test at Laboratory A and arrange for an MRI at Imaging Center B before referring the patient to a neurology colleague at Hospital C for a mental status assessment. VFL would allow for multimodal training (or inference) of a global model at the level of a central server, utilizing of these different types of data obtained from a variety different institutions, without ever having to consolidate or centralize all the information in a single data warehouse. The architecture of vertical federated learning is shown in figure 2.
Figure 2: A Vertical Federated Learning Architecture where Labratory A has protein tests, Imaging Center B has MRI Data, and Hospital C possesses psychiatric assessments. Secure entity alignment occurs before training to ensure that the patients are ”aligned” in order to perform the distributed training.
2.5 Clinical Applications for Vertical Federated Learning
VFL for disease diagnosis is far less common. Nevertheless, two notable publications are worth discussion. Firstly, Zhang and colleagues 28 focused on the security of VFL systems using an example from the Federated Technology Enabler29 (FATE). Developed by Webank’s AI department, FATE remains the world’s first industrial-grade federated learning open source framework, enabling enterprises and institutions to collaborate on data while protecting data security and privacy. Ju et al. 30 adapted the manifold projection proposed in Gao et al. 24 and applied it to federated transfer learning allowing for increased diagnostic accuracy by leveraging data diversity.
Based on the current literature review, both HFL and VFL would be ideal for DML in healthcare. Since transfer of PHI data to a central data warehouse or among collaborators never has to occur, all federated learning models inherently would be HIPAA- and HITECH-compliant; dominion over data would remain locally. The choice between and HFL and VFL model should be based on data distribution. If only a single type of data (such as MRI imaging) is going to be used for training a DML, then a HFL framework readily can be executed. Although multimodal data training can occur with HFL, either (1) the multimodal features must be present on every federate or (2) the features must project down to the same latent space, inevitably leading to a loss of information and accuracy. On the other hand, VFL could have the additional benefit of being able to handle diverse datasets. While both the security and improved accuracy of a VFL model in healthcare have been established, to date, no one has investigated how a VFL model could be trained on multimodal data.
3 Methods
3.1 Dataset and Preprocessing of the Data
The second version of Open Access Series of Imaging Studies (OASIS-2) is comprised of a longitudinal cohort of 150 individuals aged 60 to 96 years 31. All subjects were right-handed and consisted of both male and female subjects. Table 1 summarizes the list of features provided in the OASIS-2 dataset.
Table 1: Features in the OASIS-2 Dataset
Feature Description
Group Non-Demented, Demented, or Converted
Age Age at time of image acquisition (years)
Sex Sex (M or F)
EDUC Years of education
SSE Socioeconomic status as assessed by the Hollingshead Index of Social Position 32
MMSE Mini-Mental State Examination score 33
CDR Clinical Dementia Rating 34
ASF Atlas scaling factor (unitless). Computed scaling factor that transforms native-space brain and skull to the atlas target (i.e., the determinant of the transform matrix) 35
eTIV Estimated total intracranial volume (cm3) 35
nWBV Normalized whole-brain volume, expressed as a percent of all voxels in the atlas-masked image that are labeled as gray or white matter by the automated tissue segmentation process 36
The ASF, eTTV, and nWBV required MRI brain scans. Each participant underwent two or more T1-weighted MRI scanning sessions, with a minimum one-year interval in between sessions, resulting in a total of 373 imaging sessions. There are 3 to 4 individual images used from each MRI scanning session.
Among the participants, 72 individuals maintained a non-demented status throughout the study. 64 individuals were initially classified as demented during their initial assessments and retained this classification in subsequent scans, including 51 individuals with mild to moderate Alzheimer’s disease. Most notably, 14 subjects initially characterized as non-demented were subsequently diagnosed with dementia. These subjects are labeled as ”converted” on the study. The scatterplot in figure 3 illustrates the loading vectors of each demographic and clincal feature and each patient is projected down to a 2-dimensional space, using principal component analysis.
Figure 3: Scatter plot of the dataset projected onto two principal components with loading vectors
After excluding all patient records that had any missing features, the dataset comprised of 142 patients, 354 imaging sessions, and 1297 MRI images. I then normalized the values of the features from zero to one, choosing the ”Group” feature as the label. The first 102 patients comprised the training set, 20 patients as the validation set, and another 20 patients as the test set. As mentioned earlier, each patient had multiple MRI images. Therefore, our training set included 952 records, the validation set had 181 records, and the test set consisted of 164 records. Since our model cannot train on patient-level data, results are reported at the record level.
MRI images were imported in the form of NIfTI (Neuroimaging Informatics Technology Imitative) images 37 and then converted into tensors with NiBabel 38. Although a FATE compatible data loader to load the entire NiFTI images could have been used and writing a dedicated dataloader for NiFTI images would have been more useful in a clinical setting, such a feat would require approximately 1 GB of data per batch since our batch size is 32 and each NIfTI image is approximately 30 MB. Therefore, in the interest of efficiency, the MRI images were pre-processed into 256x256 JPEG format for training purposes. The 70th slice, 100th slice, and 125th slice were used to form the 3 channels of our JPEG image. Cross sectional views of a representative MRI image from our dataset is shown in figure 4.
Figure 4: Cross sectional MRI Image
3.2 Experimental Model
Image data resided on Federate A. Demographic data along with the Clinical Dementia Rating resided on Federate B. On Federate A, the bottom model used a modified ResNet 18 architecture39, a deep CNN employing residual connections to prevent vanishing gradients, architecture where the last layer of ResNet 18 was connected to a dense layer with 10 outputs. The accuracy of 3D CNNs compared to 2D CNNs on MRIs has been reported previouisly.40 Because– on average– 2D CNNs outperformed 3D counterparts, 40 I opted to use a 2D CNN. On Federate B, the bottom model was a feed-forward neural network consisting of 12 inputs (each corresponding to a demographic feature), a hidden layer with 20 neurons, and an output layer with 10 neurons. These models were subsequently interconnected through an interactive layer with a size of 10 neurons. The top layer of this VFL model also was situated on Federate B because of the label’s residence. This top layer consisted of a feed-forward neural network with 10 inputs and the resultant output. The Framework’s architecture is illustrated in figure 5.
Figure 5: A Multimodal Vertical Federated Learning Model for the Diagnosis of Alzheimer’s Dementia.
FATE allows a user to submit a task such as a training or inference job through FATE Flow. For our purpose, we used the FATE Standalone client to simulate multiple federates on a single system (FATE will still simulate all of the network connections and processes necessary). A configuration for the architecture outlined in the previous paragraph was specified. The job then was submitted to FATE flow.
Unfortunately, FATE does not support loss metrics with a validation set. In order to get insights about overfitting and underfitting, a local version of this model was designed and trained in a similar manner described above, with the exception that the interactive layer was replaced with a linear layer and all models and data resided within the same system. Hence, propagation of gradients was done locally, rather than over the network. After designing the architectural framework, the model then was executed.
4 Results
This multimodal vertical federated neural network demonstrated promising results in the detection of Alzheimer’s disease, achieving an overall accuracy rate of 82.9% on the validation dataset. This performance metric demonstrates the model’s ability to correctly classify instances of Alzheimer’s, showcasing its potential utility in clinical applications. A 3×3 confusion matrix, containing True versus Predicted classifications for three classes of test subjects is provided in Figure 6. To contextualize this accuracy rate, this model’s performance is compared with existing ligerature and benchmark models in Table 2.
Figure 6: Confusion matrix of our proposed architecture for our 164 test records.
Table 2: Comparison of results for different federated methods for AD detection
Paper Dataset Method Type Acc
Khalil et al. 18 ADNI Plasma 19 MLP HFL 89%
SVM HFL 62%
KNN HFL 67%
DT HFL 59%
LR HFL 59%
Lei et al. 21 ADNI 19, AIBL 20 FedDAvT HFL 89%
Huang et al. 22 ADNI 19, OASIS 1 23, AIBL 20 FedCM HFL N/A
OASIS 1 23 FedCM HFL 74.5%
AIBL 20 FedCM HFL 76%
Mandal (2024) OASIS 2 31 ResNet + DNN VFL 82%
Training and validation loss for the local version of this VFL architecture is illustrated in Figure 7. As explained above, local training was used to tune the hyperparameters of this federated model. With the chosen hyperparemeters, the loss plot indicates that the number of optimal epochs is 7. Training and validation loss for the local version of our architecture is illustrated in figure 7.
Figure 7: Training and Validation Loss for the local VFL Model
5 Discussion
This paper proposed and then investigated how a novel VFL platform for DML can be used for multimodal training. As discussed above, multimodal training is paramount for DML when different types of data are used for clinical diagnoses. Specific findings include the following:
Firstly, the VFL model could be trained for the diagnosis of AD, with an accuracy rate of 82.9%. This diagnostic accuracy rate is well within the results of other investigators who applied HFL for AD detection and reported accuracy rates between 62% and 89% 18; 21; 22. Despite the accuracy rate obtained, one limitation of the study is OASIS-2 dataset’s small sample size. The OASIS-2 dataset only had 150 patients, which made it challenging to train a robust neural network. Splitting the train, validation, and test sets on the imaging session or MRI image instead of on the patient axis may have yielded a higher accuracy rate. However, such a strategy also could have resulted in an over-fit because of the small sample size. Moving forward, the OASIS-3 dataset 41 consists of 1,379 patients. Additionally, more data has been included– such as both amyloid and metabolic positron emission tomography scans, neuropsychiatric testing, genetic testing, and other clinical data. Future studies, therefore, could use OASIS-3 for multimodal training of a VFL architecture for AD detection. OASIS-3 training of a VFL archtitecture most likely would produce a much more robust model trained on more modalities of data. Although training on a wide array of different datasets– as found in OASIS-3– should increase diagnostic accuracy, the availability of more data in a clinical setting could bring about further challenges if used to train the herein described VFL model. For example, the assumption that all of the demographic features rely on one federate and the image data rely on another might neither be realistic or feasible. To obviate this hurdle, one option would be to apply the Cascade Vertical Federated Learning (CVFL) framework 42 . CVFL decomposes the neural network into subnetworks, thereby allowing the utilization of partitioned labels for privacy-preserving training. CVFL allows for VFL training when some of the labels are distributed horizontally. Such a framework is advantageous particularly in scenarios where features and labels are distributed among different parties, Therefore, CVFL aligns with most clinical settings, in which data and labels often are distributed across multiple federates. The potential implementation of CVFL in future iterations of our VFL research promises a more robust and clinically relevant approach to federated learning within the healthcare domain.
Secondly, by using fully homomorphic encryption to prevent the leakage of model weights, my VFL model is privacy-preserving. Since federated learning does not require sending personal health information to the central server, it inherently is HIPAA compliant. The importance of security and privacy protections is a costly concern. With regard to preserving personal health information, in 2019, it was reported that the total cost of healthcare sector data breaches totaled \$8 million dollars, or \$400 dollars per each medical record. In the first 6 months of 2024, there already have been 340 data breaches, affecting 42,250,702 individuals 43; 44. By the end of this year, a total cost of 11 million dollars or 500 dollars per medical record could be anticipated. Thus, DML must have safeguard around PHI, and– as discussed above– all federated learning frameworks inherently would protect PHI.
Thirdly, this paper demonstrates how current, open-source federated libraries (i.e. FATE) and deep learning algorithms (i.e. ResNet-18) readily can be applied to formulate a VFL architecture for clinical diagnoses. Based on the above findings, either VFL or CVFL should be the primary focus whenever multimodal training is required in DML. These models not only work on diverse datasets simultaneously but also provide improved accuracy rates and PHI protection
In spite of the challenges and limitations outlined above, the work herein holds significant promise for further investigation of VFL implementation in DML. As reviewed above, a notable gap in the literature– regarding the use of VFL to address data segmentation challenges that arise when different features and imagery for a patient are distributed across multiple parties– currently exists. By determining whether a VFL framework for AD detection could be implemented, the challenges of data segmentation along the feature axis was successfully addressed while still protecting personal health information. Future VFL modeling frameworks should lead to further innovations in applying machine learning to clinical medicine.
6 Acknowledgements
The author of this paper would like to thank Aloke K. Mandal, M.D., Ph.D. for his clinical input and assistance in editing this manuscript.
Data was provided by OASIS: Longitudinal: Principal Investigators: D. Marcus, R, Buckner, J. Csernansky, J. Morris; P50 AG05681, P01 AG03991, P01 AG026276, R01 AG021910, P20 MH071616, U24 RR021382.

### v4

@@
 Abstract
-BACKGROUND: Segmentation of medical data, concerns about personal health information (PHI) breaches, and the direct and indirect costs of consolidating and managing such segmented date should motivate diagnostic machine learning (DML) researchers to identify privacy-preserving machine learning algorithms that can train on distributed or decentralized datasets of different modalities.
-Federated learning models provide such a decentralized machine learning framework in which multiple investigators in possession of disparate datasets and working on different devices or servers can train collaboratively a global machine learning models without ever having to exchange local data and thus can meet statutory PHI protections.
-To this end, a vertical federate learning model is devised and tested for efficacy in the detection of Alzheimer’s Disease (AD).
-METHODS: The second version of Open Access Series of Imaging Studies– with its panoply of demographic, imaging, and clinical assessment datasets– was used to test a multimodal vertical federated learning (VFL) model for AD detection.
-RESULTS: By training and validating this VFL model on the demographic, clinical, and MRI data in OASIS-2, an 82.9% accuracy rate is achieved, consistent with previously reported results.
-CONCLUSIONS: The VFL architecture proposed herein offers a novel distributed architecture, enabling collaborative learning across diverse sources of medical data while respecting statutory privacy constraints.
+In the era of rapidly advancing medical technologies, the segmentation of medical data has become inevitable, necessitating the development of privacy preserving machine learning algorithms that can train on distributed data.
+Consolidating sensitive medical data is not always an option particularly due to the stringent privacy regulations imposed by the Health Insurance Portability and Accountability Act (HIPAA).
+In this paper, I introduce a HIPAA compliant framework that can train from distributed data.
+I then propose a multimodal vertical federated model for Alzheimer’s Disease (AD) detection, a serious neurodegenerative condition that can cause dementia, severely impairing brain function and hindering simple tasks, especially without preventative care.
+This vertical federated learning (VFL) model offers a distributed architecture that enables collaborative learning across diverse sources of medical data while respecting privacy constraints imposed by HIPAA.
+The VFL architecture proposed herein offers a novel distributed architecture, enabling collaborative learning across diverse sources of medical data while respecting statutory privacy constraints.
 By leveraging multiple modalities of data, the robustness and accuracy of AD detection can be enhanced.
@@
 1 Introduction
-The increasing volume of medical data should be a priceless asset for advancing diagnostic machine learning (DML) models in healthcare 1; 2; 3.
+The increasing volume of medical data should be a priceless asset for advancing diagnostic machine learning (DML) models in healthcare [1, 2, 3].
 Commonly, these large datasets are consolidated onto a centralized server for collaborators to investigate.
 Unfortunately, consolidating medical data in such a fashion in order to train a centralized diagnostic machine learning model can be burdensome, costly, and impractical.
-For example, the Health Insurance Portability and Accountability Act (HIPAA) 4; 5 and as updated through the Health Information Technology for Economic and Clinical Health Act (HITECH)6 mandate safeguards to protect any personal health information (PHI) from any breaches.
+For example, the Health Insurance Portability and Accountability Act (HIPAA) [4, 5] and as updated through the Health Information Technology for Economic and Clinical Health Act (HITECH) [6] mandate safeguards to protect any personal health information (PHI) from any breaches.
 Additionally, consolidating medical data in order to train a centralized model often can be an impractical endeavor.
-The segmentation inherent in medical data across diverse sources imposes significant limitations on both the depth and breadth of available information for robust DML training 7.
+The segmentation inherent in medical data across diverse sources imposes significant limitations on both the depth and breadth of available information for robust DML training [7].
 Not surprisingly, consolidating and managing this data is fraught with administrative burdens, technical challenges, and considerable expenses.
-Specifically, regional health information organizations (RHIOs) require US \$12 million each for development and incur \$2-3 million in operating costs 8; 9.
+Specifically, regional health information organizations (RHIOs) require US \$12 million each for development and incur \$2-3 million in operating costs [8, 9].
 Just as a physician must grasp different types of clinical data (such as a history, physical examination, laboratory results, imaging studies, and questionnaires) to develop a unifying diagnosis, each data source used to train a DML model provides only a partial perspective.
@@
 Thus, the utility of multimodal models becomes clearly apparent.
-Arguably, as explained above, prevailing approaches to multimodal model training practice data consolidation 10 –fundamentally misaligned with the objectives of (1) preserving data segmentation, (2) protecting PHI, and (3) still optimizing model performance.
+Arguably, as explained above, prevailing approaches to multimodal model training practice data consolidation [10] –fundamentally misaligned with the objectives of (1) preserving data segmentation, (2) protecting PHI, and (3) still optimizing model performance.
 On the other hand, the integration of diverse data modalities holds promise for a more comprehensive understanding, particularly in addressing complex medical conditions.
@@
 The decision to address AD not only stems from a technical standpoint but also aligns with a broader, urgent societal need.
-Specifically, in 2024, an estimated 6.9 million or 1 in 9 (11%) of all Americans aged 65 years and older are living with AD11.
-The Centers for Medicare and Medicaid Services shall pay out US\$ 206 billion for direct AD care12.
-Highlighting the overall the overall economic and associated burdens of AD to patients, their families, and society is the fact that there is an additionally estimated 18 billion unpaid hours for dementia-related care in 2022, valued at \$339.5 billion 13.
+Specifically, in 2024, an estimated 6.9 million or 1 in 9 (11%) of all Americans aged 65 years and older are living with AD [11].
+The Centers for Medicare and Medicaid Services shall pay out US\$ 206 billion for direct AD care [12].
+Highlighting the overall the overall economic and associated burdens of AD to patients, their families, and society is the fact that there is an additionally estimated 18 billion unpaid hours for dementia-related care in 2022, valued at \$339.5 billion [13].
 The number of Americans aged 65 and older not only will increase exponentially in the United States but also elsewhere– as will the cost and other burdens of caring for AD.
@@
 2.1 What Is Federated Learning?
-Federated learning is a decentralized machine learning framework which allows multiple disparate parties– often on separate devices or servers– to collaboratively train a global machine learning model without ever having to exchange their local data 14.
+Federated learning is a decentralized machine learning framework which allows multiple disparate parties– often on separate devices or servers– to collaboratively train a global machine learning model without ever having to exchange their local data [14].
 In scenarios wherever data security, patient privacy, and maintenance of data locality are necessary, federated learning can be a powerful tool.
@@
 2.2 Horizontal Federated Learning
-Horizontal Federated Learning (HFL) is defined as a scenario in which different parties share the same feature space but contain different data samples 15.
+Horizontal Federated Learning (HFL) is defined as a scenario in which different parties share the same feature space but contain different data samples [15].
 After initialization of a global model, each client, possessing the same type of data, trains the model locally.
@@
 This averaging process ensures that the global model benefits from the diverse knowledge learned on different clients while preserving privacy.
-One common method used in this process is Federated Averaging (FedAvg) 14, where the central server orchestrates the federated learning process by specifically averaging the model updates and then redistributing the improved model to all clients.
+One common method used in this process is Federated Averaging (FedAvg) [14], where the central server orchestrates the federated learning process by specifically averaging the model updates and then redistributing the improved model to all clients.
 Therefore, HFL is effective in scenarios where multiple federates are collecting similar data samples from multiple federates (or entities) without having to consolidate this data onto a centralized server.
@@
 2.3 Clinical Applications for Horizontal Federated Learning
-There are many survey papers that both propose and implement HFL in the healthcare research 16; 17.
-In terms of primary research of HFL models for AD detection training, Khalil and colleagues 18 introduce a hardware accelerated horizontal federated multi-layer perceptron on field programmable gate arrays with 89% accuracy on blood plasma data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) 19.
+There are many survey papers that both propose and implement HFL in the healthcare research [16, 17].
+In terms of primary research of HFL models for AD detection training, Khalil and colleagues [18] introduce a hardware accelerated horizontal federated multi-layer perceptron on field programmable gate arrays with 89% accuracy on blood plasma data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) [19].
 They also used K-nearest neighbors (KNN), support vector machines, decision trees, and logistic regression analyses.
-Similarly, MRIs from the ADNI and Australian Imaging, Biomarkers and Lifestyle (AIBL) 20 datasets were utilized to implement a federated domain adaptation framework via Transformer (FedDAvT)21.
+Similarly, MRIs from the ADNI and Australian Imaging, Biomarkers and Lifestyle (AIBL) [20] datasets were utilized to implement a federated domain adaptation framework via Transformer (FedDAvT) [21].
 This approach performed Unsupervised Deep Domain Adaptation and managed to achieve the following accuracy rates of: (1) 88.75% for the AD vs.
@@
 MCI binary classification tasks.
-A process called federated conditional mutual learning (FedCM)–which considers the similarity between clients– improved these DML models that trained on MRI imaging 22.
-Using MRIs from the first version of Open Access Series of Imaging Studies (OASIS 1) 23, ADNI, and AIBL datasets and then by extending upon FedAvg through FedCM, accuracy rates of 75.5% accuracy rate on OASIS-1 and of 76.0% on AIBL and ADNI were achieved.
+A process called federated conditional mutual learning (FedCM)–which considers the similarity between clients– improved these DML models that trained on MRI imaging [22].
+Using MRIs from the first version of Open Access Series of Imaging Studies (OASIS 1) [23], ADNI, and AIBL datasets and then by extending upon FedAvg through FedCM, accuracy rates of 75.5% accuracy rate on OASIS-1 and of 76.0% on AIBL and ADNI were achieved.
 In each of the above studies of HFL models for AD detection, only one modality of data was used– whether blood plasma protein levels or MRIs.
 Nevertheless, some investigators have attempted to train HFLs on multimodal data.
-For example, Gao and colleagues performed a manifold projection of each electroencephalogram (EEG) signal, first by calculating a domain loss on each disparate signal and then by applying HFL to classify each signal24.
+For example, Gao and colleagues performed a manifold projection of each electroencephalogram (EEG) signal, first by calculating a domain loss on each disparate signal and then by applying HFL to classify each signal [24].
 Although only one type of testing was analyzed (EEGs), manifold projection is necessary since each EEG signal comes from a different area of the brain and thus can be considered multimodal data.
@@
 A special graph generative adversarial network is then used to fill out missing nodes in the graph.
-Therefore, the multimodal data is projected onto a uniform feature space, where FedAVG is applied to the locally trained graph for detecting both Alzheimer’s Disease (AD) and Autism Spectrum Disorder (ASD) 25.
-Both of the technologies described above succeeded in their objectives, training either a single federated neural network using an HFL process which would not perform well without the manifold projection 24 or on multimodal data by using the same manifold projection25.
+Therefore, the multimodal data is projected onto a uniform feature space, where FedAVG is applied to the locally trained graph for detecting both Alzheimer’s Disease (AD) and Autism Spectrum Disorder (ASD) [25].
+Both of the technologies described above succeeded in their objectives, training either a single federated neural network using an HFL process which would not perform well without the manifold projection [24] or on multimodal data by using the same manifold projection [25].
 Unfortunately, projecting features onto a different latent space inevitably results in loss of information.
@@
 2.4 Vertical Federated Learning
-Vertical Federated Learning (VFL)– in contrast to HFL– addresses instances in which multiple federates share the feature space and each training example is spread across different entities 26.
+Vertical Federated Learning (VFL)– in contrast to HFL– addresses instances in which multiple federates share the feature space and each training example is spread across different entities [26].
 Clients possess different features of a common entity, and the goal is to train a model that leverages all features for improved prediction and analysis.
-This approach enables predictive modeling and analytics across heterogeneous, multimodal data sources while preserving data privacy and security 27.
+This approach enables predictive modeling and analytics across heterogeneous, multimodal data sources while preserving data privacy and security [27].
 In the clinical setting, for example, if a primary care physician (PCP) is concerned that a patient has early signs of AD, the PCP might order a test at Laboratory A and arrange for an MRI at Imaging Center B before referring the patient to a neurology colleague at Hospital C for a mental status assessment.
@@
 Nevertheless, two notable publications are worth discussion.
-Firstly, Zhang and colleagues 28 focused on the security of VFL systems using an example from the Federated Technology Enabler29 (FATE).
+Firstly, Zhang and colleagues [28] focused on the security of VFL systems using an example from the Federated Technology Enabler (FATE) [29].
 Developed by Webank’s AI department, FATE remains the world’s first industrial-grade federated learning open source framework, enabling enterprises and institutions to collaborate on data while protecting data security and privacy.
 Ju et al.
-30 adapted the manifold projection proposed in Gao et al.
-24 and applied it to federated transfer learning allowing for increased diagnostic accuracy by leveraging data diversity.
+[30] adapted the manifold projection proposed in Gao et al.
+[24] and applied it to federated transfer learning allowing for increased diagnostic accuracy by leveraging data diversity.
 Based on the current literature review, both HFL and VFL would be ideal for DML in healthcare.
@@
 3.1 Dataset and Preprocessing of the Data
-The second version of Open Access Series of Imaging Studies (OASIS-2) is comprised of a longitudinal cohort of 150 individuals aged 60 to 96 years 31.
+The second version of Open Access Series of Imaging Studies (OASIS-2) is comprised of a longitudinal cohort of 150 individuals aged 60 to 96 years [31].
 All subjects were right-handed and consisted of both male and female subjects.
@@
 EDUC Years of education
-SSE Socioeconomic status as assessed by the Hollingshead Index of Social Position 32
-MMSE Mini-Mental State Examination score 33
-CDR Clinical Dementia Rating 34
-ASF Atlas scaling factor (unitless). Computed scaling factor that transforms native-space brain and skull to the atlas target (i.e., the determinant of the transform matrix) 35
-eTIV Estimated total intracranial volume (cm3) 35
-nWBV Normalized whole-brain volume, expressed as a percent of all voxels in the atlas-masked image that are labeled as gray or white matter by the automated tissue segmentation process 36
+SSE Socioeconomic status as assessed by the Hollingshead Index of Social Position [32]
+MMSE Mini-Mental State Examination score [33]
+CDR Clinical Dementia Rating [34]
+ASF Atlas scaling factor (unitless). Computed scaling factor that transforms native-space brain and skull to the atlas target (i.e., the determinant of the transform matrix) [35]
+eTIV Estimated total intracranial volume (cm3) [35]
+nWBV Normalized whole-brain volume, expressed as a percent of all voxels in the atlas-masked image that are labeled as gray or white matter by the automated tissue segmentation process [36]
 The ASF, eTTV, and nWBV required MRI brain scans.
@@
 Since our model cannot train on patient-level data, results are reported at the record level.
-MRI images were imported in the form of NIfTI (Neuroimaging Informatics Technology Imitative) images 37 and then converted into tensors with NiBabel 38.
+MRI images were imported in the form of NIfTI (Neuroimaging Informatics Technology Imitative) images [37] and then converted into tensors with NiBabel [38].
 Although a FATE compatible data loader to load the entire NiFTI images could have been used and writing a dedicated dataloader for NiFTI images would have been more useful in a clinical setting, such a feat would require approximately 1 GB of data per batch since our batch size is 32 and each NIfTI image is approximately 30 MB.
@@
 Demographic data along with the Clinical Dementia Rating resided on Federate B.
-On Federate A, the bottom model used a modified ResNet 18 architecture39, a deep CNN employing residual connections to prevent vanishing gradients, architecture where the last layer of ResNet 18 was connected to a dense layer with 10 outputs.
-The accuracy of 3D CNNs compared to 2D CNNs on MRIs has been reported previouisly.40 Because– on average– 2D CNNs outperformed 3D counterparts, 40 I opted to use a 2D CNN.
+On Federate A, the bottom model used a modified ResNet 18 architecture [39], a deep CNN employing residual connections to prevent vanishing gradients, architecture where the last layer of ResNet 18 was connected to a dense layer with 10 outputs.
+The accuracy of 3D CNNs compared to 2D CNNs on MRIs has been reported previouisly [40].
+Because– on average– 2D CNNs outperformed 3D counterparts [40], I opted to use a 2D CNN.
 On Federate B, the bottom model was a feed-forward neural network consisting of 12 inputs (each corresponding to a demographic feature), a hidden layer with 20 neurons, and an output layer with 10 neurons.
@@
 Paper Dataset Method Type Acc
-Khalil et al. 18 ADNI Plasma 19 MLP HFL 89%
+Khalil et al. [18] ADNI Plasma [19] MLP HFL 89%
 SVM HFL 62%
@@
 LR HFL 59%
-Lei et al. 21 ADNI 19, AIBL 20 FedDAvT HFL 89%
-Huang et al. 22 ADNI 19, OASIS 1 23, AIBL 20 FedCM HFL N/A
-OASIS 1 23 FedCM HFL 74.5%
-AIBL 20 FedCM HFL 76%
-Mandal (2024) OASIS 2 31 ResNet + DNN VFL 82%
+Lei et al. [21] ADNI [19], AIBL [20] FedDAvT HFL 89%
+Huang et al. [22] ADNI [19], OASIS 1 [23], AIBL [20] FedCM HFL N/A
+OASIS 1 [23] FedCM HFL 74.5%
+AIBL [20] FedCM HFL 76%
+Mandal (2024) OASIS 2 [31] ResNet + DNN VFL 82%
 Training and validation loss for the local version of this VFL architecture is illustrated in Figure 7.
@@
 Firstly, the VFL model could be trained for the diagnosis of AD, with an accuracy rate of 82.9%.
-This diagnostic accuracy rate is well within the results of other investigators who applied HFL for AD detection and reported accuracy rates between 62% and 89% 18; 21; 22.
+This diagnostic accuracy rate is well within the results of other investigators who applied HFL for AD detection and reported accuracy rates between 62% and 89% [18, 21, 22].
 Despite the accuracy rate obtained, one limitation of the study is OASIS-2 dataset’s small sample size.
@@
 However, such a strategy also could have resulted in an over-fit because of the small sample size.
-Moving forward, the OASIS-3 dataset 41 consists of 1,379 patients.
+Moving forward, the OASIS-3 dataset [41] consists of 1,379 patients.
 Additionally, more data has been included– such as both amyloid and metabolic positron emission tomography scans, neuropsychiatric testing, genetic testing, and other clinical data.
@@
 For example, the assumption that all of the demographic features rely on one federate and the image data rely on another might neither be realistic or feasible.
-To obviate this hurdle, one option would be to apply the Cascade Vertical Federated Learning (CVFL) framework 42 .
+To obviate this hurdle, one option would be to apply the Cascade Vertical Federated Learning (CVFL) framework [42] .
 CVFL decomposes the neural network into subnetworks, thereby allowing the utilization of partitioned labels for privacy-preserving training.
@@
 With regard to preserving personal health information, in 2019, it was reported that the total cost of healthcare sector data breaches totaled \$8 million dollars, or \$400 dollars per each medical record.
-In the first 6 months of 2024, there already have been 340 data breaches, affecting 42,250,702 individuals 43; 44.
+In the first 6 months of 2024, there already have been 340 data breaches, affecting 42,250,702 individuals [43, 44].
 By the end of this year, a total cost of 11 million dollars or 500 dollars per medical record could be anticipated.